An ac generator has a frequency of 7.9 kHz and a voltage of 41 V. When an inductor is connected between the terminals of this generator, the current in the inductor is 47 mA. What is the inductance of the inductor?
L =   
Answer #1

frequency , f = 7.9 x 103 Hz

Voltage, V = 41 volts

Current , i = 47 mA = 47 x 10-3 A

Let the inductance be ''L''

then current in inductor is given as ::

i= V/2\pifL

L = V/2\piif

inserting the above values

L = 41 / 2 (3.14) (7.9 x 103 )(47 x 10-3)

L = 0.0176 H
Answer #2

Here is what I solved before, please modify the figures as per your question. Please let me know if you have further questions. Ifthis helps then kindly rate 5-stars.

An ac generator has a frequency of 7.0 kHz and a voltage of 30 V. When an inductor is connected between the terminals of this generator, the current in the inductor is 65 mA. What is the inductance of the inductor?

Answer

Given

Frequency of AC generator f = 7000 Hz

   voltage    V = 30 V

    Current   I   = 0.065 A

   inductive reactance   

           X_L =    V / I

                  =   30 / 0.065

                  = 461.538

          2 ? fL = 461.538

     inductance of the inductor

                     L =    461.538 / ( 2*3.14 * 7000 )

                        = 0.01 H